OpenStudy (anonymous):

(2x+5)/(x^3-1) Find the limit of x as x approaches infinity Must use only algebraic methods...

OpenStudy (anonymous):

devide everything by x and take the limit, keeping in mind that \(a/\infty=0\), where a is any number

OpenStudy (anonymous):

Im sorry but I cant seem to figure out how to write the equation as divided by x. Sorry for my inability to preform simple math ;)

OpenStudy (anonymous):

Mate?

OpenStudy (anonymous):

I'm pretty sure this is the same problem as the last one. I did if algebraically after. but here it is in one step. \[\frac{ 2x+5 }{ x^3 - 1 } = \frac{ 2+ \frac{ 5 }{ x } }{ x^2 - \frac{ 1 }{ x } }\]

OpenStudy (anonymous):

Haha it was, but I couldn't seem to decode the algebraic method you used to work on other problems so I had to ask it again

OpenStudy (anonymous):

hehe. do you see it now? every term is simply divided individually by x

OpenStudy (anonymous):

Well when you put it that way, it makes me seem like an idiot; So I am going to just say I knew that originally and that I didn't just realize that....

OpenStudy (anonymous):

lol :P

OpenStudy (anonymous):

So thanks!...again...

OpenStudy (anonymous):

glad i coudl help